### v1

# Implementation of Time-Varying Controllers for a Nonholonomic Mobile Robot: Experimental Studies

Abstract
We consider a kinematic model of a wheeled mobile robot controlled by translational and angular velocities. For this class of nonholonomic systems, a family of time-varying feedback controllers was proposed in our previous works using gradient flow approximation techniques. In the present study, these controllers are implemented on a TurtleBot3 Burger (TB3) mobile robot to provide experimental validation of the stabilization problem with oscillating input signals. In addition, the admissibility problem of a gradient flow is investigated to justify the construction of a Lyapunov function candidate. The presented experimental results demonstrate the possibility of stabilizing the reference position of the robot using feedback controls with practically acceptable parameters.
keywords:
nonholonomic system, stabilization, Lyapunov function, time-varying feedback, sampling, mobile robot
1 Introduction
Control design problems for nonholonomic systems have attracted considerable interest from analytical [4, 16, 11, 6, 14], computationally-oriented [10, 15, 13, 5], and data-driven [12, 9] perspectives. A unified approach to solving the stabilization, motion planning, and obstacle avoidance problems was proposed in [8], based on approximating the gradient flow of a certain potential function. Such a potential function plays the role of a Lyapunov function candidate in the stabilization problem. Although it was proven that this strategy guarantees convergence of trajectories to the steady state for any positive-definite Lyapunov function candidate satisfying additional regularity conditions, the problem of optimizing the transient behavior remains open. In particular, numerical simulations show chattering phenomena when sampling-based control strategies are implemented (cf. [7]). This chattering arises from the oscillatory nature of these time-varying controllers. In a realistic setup with physical limitations, the input signals vary within velocity constraints, causing considerable chattering around non-admissible directions. As a result, the control system is forced to switch repeatedly, leading to zigzagging behavior.
In this work, we consider a variational formulation aimed at minimizing chattering through a subsequent approximation of gradient-based control techniques. The proposed approach is based on measuring the admissibility of a given gradient flow generated by a potential function when applied to a nonholonomic system over a specified domain of the state space. The admissibility measure of the resulting set of trajectories, or phase portrait, within this domain is characterized in terms of an integral measure that accounts for the magnitude of the gradient of the generating potential function.
Note that the convergence proofs previously presented in [8] require a certain design parameter $\varepsilon>0$ to be sufficiently small, which leads to high frequencies of order $1/\varepsilon$ in the oscillatory components of the controllers. However, implementing such controllers in real nonholonomic systems requires constraining the frequency within a given bandwidth while simultaneously satisfying physical control constraints. The present paper therefore aims to provide a proof of concept for the practical applicability of oscillatory controllers that account for acceptable ranges of frequency parameters and control gains.
2 Stabilizing Controls for the Unicycle Kinematics
In this paper, we consider the well-known kinematic model of a unicycle rolling on the $(x_{1},x_{2})$-plane, described by the following system of differential equations (see, e.g., [4]):
$\dot{x}=u_{1}f_{1}(x)+u_{2}f_{2}(x),\quad x=(x_{1},x_{2},x_{3})^{\top}\in{\mathbb{R}}^{3},\;u=(u_{1},u_{2})^{\top}\in{\mathbb{R}}^{2},$ (1)
where the vector fields $f_{1},f_{2}:{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ are defined as $f_{1}(x)=\big(\cos x_{3},\sin x_{3},0\big)^{\top}$, $f_{2}(x)=\big(0,0,1\big)^{\top}$. Here, $(x_{1},x_{2})$ are the coordinates of the contact point of the wheel, and $x_{3}$ is the angle between the plane of the wheel and the $x_{1}$-axis. The controls $u_{1}$ and $u_{2}$ describe the translational and angular velocities of the unicycle, respectively. To simplify the presentation, all variables of this control system are assumed to be dimensionless. For the purpose of control design, we introduce the matrix ${\cal F}(x)$ composed of the vector fields $f_{1}(x)$, $f_{2}(x)$ and their Lie bracket $f_{1},f_{2}](x)=\frac{\partial f_{2}(x)}{\partial x}f_{1}(x)-\frac{\partial f_{1}(x)}{\partial x}f_{2}(x)$: F(x) = (f_1(x),f_2(x),[f_1,f_2](x)) = (cosx30 sinx3sinx30 - cosx30 1 0 ). This matrix is nonsingular for all $x\in{\mathbb{R}}^{3}$, and its inverse is given by
${\cal F}^{-1}(x)=\begin{pmatrix}\cos x_{3}&\sin x_{3}&0\\ 0&0&1\\ \sin x_{3}&-\cos x_{3}&0\end{pmatrix}$. (2)
The nonsingularity of ${\cal F}(x)$ implies that the driftless control-affine system (1) satisfies the Lie Algebra Rank Condition (LARC): span (f_1(x),f_2(x),[f_1,f_2](x)) = R^3 for all   x∈R^3. Hence, by the Chow–Rashevsky theorem, system (1) is completely controllable in ${\mathbb{R}}^{3}$. For a given real-valued potential function $V\in C^{2}({\mathbb{R}}^{3})$, the approximate motion of system (1) in the direction of the negative gradient of $V(x)$ is generated by control functions $u^{\omega}(x,t)=(u_{1}^{\omega}(x,t),u_{2}^{\omega}(x,t))^{\top}$ of the following form [8]:
$u_{1}^{\omega}(x,t)=a_{1}(x)+k_{1}\sqrt{\omega|a_{12}(x)|}\,\mathrm{sign}(a_{12}(x))\cos(\omega t),$ (3)
$u_{2}^{\omega}(x,t)=a_{2}(x)+k_{2}\sqrt{\omega|a_{12}(x)|}\sin(\omega t),$
where $\omega{=}\frac{2\pi}{\varepsilon}$ is the control frequency with sufficiently small $\varepsilon>0$, the coefficients $k_{1}$, $k_{2}$ satisfy $k_{1}k_{2}{=}4$, and the functions $a_{1}(x)$, $a_{2}(x)$, $a_{12}(x)$ are defined as
$a(x)=\left(a_{1}(x),a_{2}(x),a_{12}(x)\right)^{\top}=-\gamma\mathcal{F}^{-1}(x)\nabla V(x)^{\top},$ (4)
where $\gamma>0$ is a gain parameter, $\mathcal{F}^{-1}(x)$ is given by (2), and $\nabla V(x)=\left(\frac{\partial V(x)}{\partial x_{1}},\frac{\partial V(x)}{\partial x_{2}},\frac{\partial V(x)}{\partial x_{3}}\right)$ denotes the gradient of $V(x)$.
In [8], the particular choice $k_{1}=k_{2}=2$ was adopted. Here, we introduce distinct coefficients to provide more flexibility in tuning the control parameters. In particular, as will be shown in Section 4, an appropriate selection of $k_{1}$ and $k_{2}$ allows adjustment of the amplitudes of the oscillatory inputs so that the translational and angular velocities remain within the admissible control bounds. The formulas for the control components in (4) can be written explicitly using (2):
$a_{1}(x)=-\gamma\left(\frac{\partial V(x)}{\partial x_{1}}\cos x_{3}+\frac{\partial V(x)}{\partial x_{2}}\sin x_{3}\right),\,a_{2}(x)=-\gamma\frac{\partial V(x)}{\partial x_{3}},$ (5)
$a_{12}(x)=-\gamma\left(\frac{\partial V(x)}{\partial x_{1}}\sin x_{3}-\frac{\partial V(x)}{\partial x_{2}}\cos x_{3}\right).$
As discussed in [17], there are two approaches to implementing the controls (3). The first approach is based on the sampling concept and assumes that the control amplitude vector $a(x)$ is updated at each time instant $t_{j}=\varepsilon j$ for $j=0,1,$... . In this case, the solutions of the resulting closed-loop system are treated according to the $\pi_{\varepsilon}$-definition [16, 8] with the control:
$u=u^{\omega}(x(t_{j}),t)\quad\text{for}\;\;t\in[j\varepsilon,(j+1)\varepsilon),\;\;j=0,1,\dots\;.$ (6)
For a class of positive definite functions $V$, convergence of the $\pi_{\varepsilon}$-solutions to the zero equilibrium of systems of the form (1) has been proved in [16, 8] under certain technical assumptions.
The second approach assumes that the control amplitudes evolve continuously in time, that is,
$u=u^{\omega}(x(t),t),\quad t\geq 0,$ (7)
for system (1). In the sequel, we will refer to the solution definitions corresponding to (6) and (7) as the “sampling” and “continuous-time” approaches, respectively. In Section 4, both approaches will be implemented experimentally.
3 Justification for the Lyapunov Function Candidate
The method of gradient flow approximation, developed in [8], establishes the solvability of the stabilization problem for a broad class of nonholonomic systems, including the unicycle model (1), with any sufficiently regular positive definite potential function. However, the problem of the optimal choice of such a potential function has not been addressed so far. To formulate this type of problem, we assume that the state $x$ and the control $u$ of system (1) are constrained to closed domains $X\subset{\mathbb{R}}^{3}$ and $U\subset{\mathbb{R}}^{2}$, respectively, and that $(x,u)=(0,0)$ is an interior point of $D\times U$.
For a function $V:X\to\mathbb{R}$ of class $C^{2}$, consider the gradient flow system
$\dot{\bar{x}}(t)=-\nabla V(\bar{x}(t))^{\top},\quad x(t)\in X,$ (8)
where the transpose on the right-hand side is taken under the convention that $\nabla V(\bar{x})$ is a row vector. System (8) possesses (local) existence and uniqueness of solutions to the Cauchy problem due to the assumption $V\in C^{2}(X;{\mathbb{R}})$. We call the gradient flow of $V$ admissible for system (1) in $X$ if, for each solution $\bar{x}(t)\in X$ of (8) defined on $t\in[0,+\infty)$, there exists a control $\bar{u}:[0,+\infty)\to U$ such that $x=\bar{x}(t)$ satisfies (1) with $u=\bar{u}(t)$ for all $t\in[0,+\infty)$. We then introduce the following “measure of admissibility” of the gradient flow of $V$ for system (1) in $X$ with controls in $U$:
$J_{X}[V]=\frac{1}{\mu(X)}\int_{X}\frac{\inf_{u\in U}\|u_{1}f_{1}(x)+u_{2}f_{2}(x)+\nabla V(x)\|^{q}}{\|\nabla V(x)\|^{q}}dx.$ (9)
Here, $\|\cdot\|$ denotes the Euclidean norm in $\mathbb{R}^{n}$, $\mu(X)$ is the Lebesgue measure of $X$, and $q$ is a positive parameter. Clearly, if the gradient flow of $V$ is admissible for system (1) and $\nabla V(x)\neq 0$ for almost all $x\in X$, then the integral in (9) is well-defined and $J_{X}[V]=0$. Hence, the minimization of $J_{X}$ within some class of potential functions can be considered as a receipt for improving the performance of a gradient-based control algorithm to ensure better tracking properties of the reference dynamics (8) by the closed-loop trajectories.
The numerator in the integrand of (9) is computed for system (1) by the following technical lemma.
Lemma 3.1.
For $x,p\in{\mathbb{R}}^{3}$, define
$\rho(x,p):=\inf_{u\in{\mathbb{R}}^{2}}\|u_{1}f_{1}(x)+u_{2}f_{2}(x)+p\|.$ (10)
Then
$\rho(x,p)=|p_{1}\sin x_{3}-p_{2}\cos x_{3}|\quad\text{for all}\;\;x,p\in{\mathbb{R}}^{3}.$ (11)
Proof 3.2.
The function $\rho(x,p)$, defined by (10), is the length of the projection of the vector $p\in{\mathbb{R}}^{3}$ onto the subspace orthogonal to ${\rm span}\{f_{1}(x),f_{2}(x)\}$. Thus,
$\rho(x,p)=\frac{|\left<p,f_{1}(x)\times f_{2}(x)\right>|}{\|f_{1}(x)\times f_{2}(x)\|},$ (12)
where $\left$$\cdot$$\cdot\right$> and $f_{1}(x)\times f_{2}(x)$ denote the inner product and the cross product of vectors in ${\mathbb{R}}^{3}$, respectively. Straightforward computations show that $f_{1}(x)\times f_{2}(x)=(\sin x_{3},-\cos x_{3},0)^{\top},\;\|f_{1}(x)\times f_{2}(x)\|=1,$ so (11) follows from (12).
A natural way of defining an optimal Lyapunov function candidate for system (1) in $X$ can be as follows: find a function $V\in C^{2}(X;{\mathbb{R}})$ that minimizes the cost (9) under the constraint that $V$ is positive definite in $X$. This optimization problem is highly nonlinear and nonconvex, and it poses significant analytical and computational challenges. In particular, the set of positive definite functions is nonconvex in any reasonable function space, and the functional (9) exhibits singular behavior at critical points of $V$ due to the presence of $\nabla V$. Note that, in the case $U={\mathbb{R}}^{2}$, the functional $J_{X}[V]$ is invariant under multiplication of $V$ by a nonzero constant. Therefore, a solution to the above optimization problem (if it exists) is not unique without additional constraints. A comprehensive analysis of this general problem is beyond the scope of the present paper. Instead, we restrict our attention to estimating $J_{X}[V]$ within some positive definite quadratic forms.
$c_{1}$ $c_{2}$ $c_{3}$ $J_{X}[V]$
1 1 1 0.3333
2 1 1 0.3056
0.5 1 1 0.3658
1 2 1 0.4716
1 0.5 1 0.2123
1 1 2 0.2228
1 1 0.5 0.4219
Table 1: Values of $J_{X}[V]$.
To estimate possible improvements of the admissibility measure $J_{X}[V]$ in comparison with the standard Lyapunov function candidate $V_{1}(x)=\|x\|^{2}$, we assign different scalings to the three state variables of system (1). Thus, we compute the functional $J_{X}[V]$ for the quadratic form $V(x)=c_{1}x_{1}^{2}+c_{2}x_{2}^{2}+c_{3}x_{3}^{2}$ under different choices of coefficients $c_{i}>0$. The computational results are summarized in Table 1 for the case $X=[-1,1]^{3}$, $U={\mathbb{R}}^{2}$, and $q=2$. In Table 1, the rows corresponding to parameters $(c_{1},c_{2},c_{3})$ for which $V$ ensures better performance than $V_{1}$, in the sense that $J_{X}[V]<J_{X}[V_{1}]\approx 0.3333$, are highlighted in bold. The data indicate that performance improves when $c_{1}$ and $c_{3}$ are increased and $c_{2}$ is decreased relative to the initial choice $c_{1}=c_{2}=c_{3}=1$. This observation motivates introducing the following Lyapunov function candidate for system (1):
$V_{\alpha}(x)=\alpha(x_{1}^{2}+x_{3}^{2})+\frac{x_{2}^{2}}{\alpha},\quad\alpha>1,$ (13)
where $\alpha$ is a parameter. In particular, computing the cost (9) for $V_{\alpha}$ with $\alpha=2,4,10$, we obtain: J_X[V_2]≈0.1403,  J_X[V_4]≈0.0962,  J_X[V_10]≈0.0906. The obtained sequence of decreasing cost values justifies the use of the Lyapunov function candidates (13) for further experiments. It should be emphasized that quadratic forms cannot serve as control-Lyapunov functions for systems of the form (1) due to Brockett’s necessary stabilizability condition. Indeed, Brockett’s condition implies that the considered class of nonholonomic systems does not admit any continuously differentiable control-Lyapunov function.
4 Experimental Implementation
In this section, we present the experimental results obtained by implementing the proposed theoretical framework on a real robotic platform. We evaluate two Lyapunov functions: the classical squared norm and the quadratic form (13). In addition, we compare two control approaches (6) and (7).
4.1 Experimental Setup
Robotic Platform. The experiments were conducted using a TurtleBot3 Burger (TB3) – an affordable, differential-drive mobile robot (see Figure 1, left) [1]. The robot’s control and communication were handled by its onboard Raspberry Pi 4B, running Ubuntu Server 22.04.5 LTS (64-bit) and ROS 2 Humble. Mechanically, the TB3 features two off-centered, independently actuated wheels and a caster ball, thus the robot is not perfectly unicycle. However, a differential-drive robot can be modeled as unicycle by mapping the velocities of the left and right wheel $v_{l}$ and $v_{r}$ (linear velocities at the wheel edge) to the unicycle velocities $u_{1}$ and $u_{2}$ mentioned in (1). Let $d$ be the distance between the two wheels of the differential drive robot, and assume the robot’s center of rotation is at the midpoint between the wheels. Then, the unicycle’s translational and angular velocities, respectively $u_{1}$ and $u_{2}$, are given by $u_{1}=\frac{v_{r}+v_{l}}{2},\quad u_{2}=\frac{v_{r}-v_{l}}{d}.$ The inverse of this transformation is handled automatically by the diff_drive_controller of ROS 2. It accepts the unicycle-style Twist commands on /cmd_vel and computes the corresponding wheel velocities. The robot’s motion is constrained by its kinematics with the maximum translational velocity of 0.22 m/s and the maximum angular velocity of 2.84 rad/s. No onboard sensors were used for state estimation or control, but four reflective markers were mounted for the motion capture system.
Motion Capture System. Robot pose information was obtained using an OptiTrack motion capture system, consisting of eight infrared cameras and the Motive tracking software. As previously stated, four reflective markers were attached to the top of the TB3, allowing the estimation of its x- and y-position as well as the yaw orientation. The Motive software typically computes the center point of these markers and uses this as a reference pose. However, this point did not align precisely with the robot’s pivot point. To ensure accurate control, the reference pose was adjusted manually to match the true center of rotation (see Figure 1, right). We used a dedicated wireless access point to connect to the OptiTrack system via LAN, ensuring the TB3 and a remote laptop used for robot control could communicate in a known low-latency wireless network.
Figure 1: Left: TurtleBot3 Burger; Right: TB3 marker configuration
Software Architecture. The control logic for both simulation and real-world experiments was implemented as a single ROS 2 node using the rclpy Python library. In real-world experiments, the controller subscribes to the robot pose data provided by the OptiTrack system via the NatNet protocol, whereas in simulations it subscribes to the standard /odom ROS 2 topic published by Gazebo. Control commands are published on the standard /cmd_vel topic, recognized by both the real-world TB3 and the Gazebo simulation. This allows us to use one and the same controller logic in different domains. Control commands are recomputed and sent as fast as possible which turned out to be every 0.5 milliseconds. Experimental data and controller parameters were recorded automatically for each trial. Plots were generated using Python’s matplotlib library, and all data were exported in .csv format for further analysis.
Simulation Environment. Simulation experiments were conducted using the predefined TB3 Gazebo environment provided in the Robotis e-Manual [2]. An empty world was used to replicate a simple, obstacle-free arena. The same TB3 model (Burger) as in real world was used to ensure consistency in dimensions, kinematics and dynamics. The model incorporates basic rigid body physics including link masses and joint properties allowing the robot to respond realistically to velocity commands although small inaccuracies in terms of friction and inertia compared to the real-world robot may exist. Due to the software architecture the same Python controller could be used.
Figure 2: Experiment 1 – Parameter set $(P1)$: continuous-time vs. sampling.
4.2 Experimental Results
For all experiments, the initial state is set to $x^{0}=(-0.5,-0.5,0)^{\top}$, and the target state is $x^{*}=(0,0,0)^{\top}$. The robot stops automatically if it is in the vicinity of that goal point to obtain comparable convergence time results. The translational and angular velocities are governed by the control law (3), (5) with the frequency parameter $\omega=2\pi$ (i.e., $\varepsilon=1$) and the gain $\gamma=0.05$. It is important to note that theoretical analyses typically assume sufficiently high control frequencies. However, the robot’s physical constraints prevent operation at large frequency ranges. So, to adjust control bounds to the admissible range $|u_{1}(t)|\leq 0.22$, $|u_{2}(t)|\leq 2.84$, we select relatively small values of $\omega$ and $\gamma$. Moreover, since larger admissible values are allowed for $u_{2}$, we choose $k_{1}<k_{2}$. This choice avoids the need to further reduce $\omega$ and $\gamma$ while still satisfying the input constraints.
For the experimental validation, we use the Lyapunov function candidate (13) and controls of the form (3), (5) with one of the following sets of parameters:
$\alpha$ $=1,\;k_{1}=0.5,\;k_{2}=8$ (P1)
$\alpha$ $=1,\;k_{1}={1}/{\sqrt{2}},\;k_{2}=4\sqrt{2}$ (P2)
$\alpha$ $=4,\;k_{1}=0.5,\;k_{2}=8$ (P3)
$\alpha$ $=10,\;k_{1}=0.5,\;k_{2}=8$ (P4)
In Experiments 1 and 2, we examine two different parameter selections $(P1)$ and $(P2)$ with the standard sum-of-squares Lyapunov function candidate $V_{1}(x)=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}$. By comparing the trajectories plots in these two experiment in Figs. 2 and 3, we observe that the first choice of control coefficients $k_{1}=0.5$ and $k_{2}=8$ demonstrates a better transient performance, so this choice is used in subsequent experiments.
We also compare two different approaches for defining the closed-loop solutions in each of the experiments. For the continuous-time approach (7), the control amplitude vector $a(x)$ is recomputed at each new computation of the controls $u_{1}$ and $u_{2}$. For the sampling approach (6), the coefficients are only updated after a period of $\varepsilon=1$.
In Figs. 2–5, the units of $x_{1}$ and $x_{2}$ are $m$, $x_{3}$ is in radrad, the translational velocity control $u_{1}$ is in $m/s$, and the angular velocity control $u_{2}$ is in $rad/s$. Videos of all the experiments are available on YouTube [3].
Figure 3: Experiment 2 – Parameter set $(P2)$: continuous-time vs. sampling.
Comparing the continuous-time and sampling approach, we can see that they generate clearly distinct trajectories. The continuous time approach follows a relatively direct and steep path which can especially seen from the parameter set $(P1)$, where the robot moves straight $\pi/4$ to the goal without chattering from about $1/3$ of the distance. The sampling approach produces a more curved and more chattering but faster path. When comparing the same approach across parameter sets $(P1)$ and $(P2)$, the overall trajectory shape is preserved, but parameter set $(P2)$ exhibits higher chattering for both approaches, thus $k_{1}$ and $k_{2}$ mostly influence chattering. Across the parameter sets, the two approaches maintain their characteristic differences. Both continuous and sampling approaches have a higher convergence time for parameter set $(P2)$ due to the increased chattering whereas this is more visible for the continuous time approach as the robot was almost not moving in a straight line to the target.
Figure 4: Experiment 3 – Parameter set $(P3)$: continuous-time vs. sampling.
Figure 5: Experiment 4 – Parameter set $(P4)$: continuous-time vs. sampling.
Varying the $\alpha$-parameter in the Lyapunov function candidate (13) drastically changed the shape of the robot’s trajectory. In Figs. 4 and 5, we observe a very fast convergence in $x_{1}$, whereas the overall convergence time increased as the robot had then to move in the non-admissible direction $x_{2}$, i.e. sideways. Across both parameter sets $(P3)$ and $(P4)$, the sampling approach is slower than the continuous-time one.
We observe in Figs. 2–5 that the control signals $u_{1}$ and $u_{2}$, evaluated along the presented trajectories, perfectly satisfy the control constraints specified above, and the transient behavior is acceptable for practical applications.
5 Conclusions and Future Work
The main contribution of this work is twofold. First, the possibility of practical application of the gradient-based controllers for tracking a real mobile robot toward a prescribed target position is justified. It is remarkable to note that although the theoretical stability proof is valid for sufficiently small values of the parameter $\varepsilon$ (and therefore for sufficiently large control frequency $\omega$), our experiments demonstrate the applicability of these controllers at moderate frequencies with $\varepsilon=1$. The choice of control coefficients $k_{1}$ and $k_{2}$ is justified in the present work based on control constraints for differential-drive robots combined with gradient-based control formulas under a realistic choice of a Lyapunov function candidate.
Second, through the formulation of a suitable optimization problem, an appropriate class of Lyapunov candidates has been studied. The construction of the first terms of a minimizing sequence demonstrates the potential of this approach to reduce chattering behavior and thus minimize zigzagging in the closed-loop dynamics. These observations are well confirmed in Experiments 3 and 4, where the initial stage of the closed-loop trajectory exhibits significantly more regular behavior compared to the strongly zigzagging evolution obtained with a standard sum-of-squares Lyapunov function candidate.
A further research direction could focus on studying the well-posedness of the associated optimization problems and on developing efficient computational schemes for constructing Lyapunov function candidates. We leave this open theoretical question for future investigation.
Another promising direction concerns experimental research aimed at achieving guaranteed finite-time practical stability. In this context, by exploiting non-differentiable Lyapunov function candidates, we expect to characterize the time required to reach a prescribed neighborhood of the target point using gradient-flow-type systems. This experimental study is left for future development, possibly involving non-Lipschitz generating potential functions.
Acknowledgement
Alexander Zuyev gratefully acknowledges the funding by the European Regional Development Fund (ERDF) within the programme Research and Innovation — Grant Number ZS/2023/12/182138.

### v2

@@
 2 Stabilizing Controls for the Unicycle Kinematics
-In this paper, we consider the well-known kinematic model of a unicycle rolling on the $(x_{1},x_{2})$-plane, described by the following system of differential equations (see, e.g., [4]):
+In this paper, we consider the well-known kinematic model of a unicycle rolling on the $(x_{1},x_{2})$-plane, described by the following system (see, e.g., [5]):
 $\dot{x}=u_{1}f_{1}(x)+u_{2}f_{2}(x),\quad x=(x_{1},x_{2},x_{3})^{\top}\in{\mathbb{R}}^{3},\;u=(u_{1},u_{2})^{\top}\in{\mathbb{R}}^{2},$ (1)
@@
 To simplify the presentation, all variables of this control system are assumed to be dimensionless.
-For the purpose of control design, we introduce the matrix ${\cal F}(x)$ composed of the vector fields $f_{1}(x)$, $f_{2}(x)$ and their Lie bracket $f_{1},f_{2}](x)=\frac{\partial f_{2}(x)}{\partial x}f_{1}(x)-\frac{\partial f_{1}(x)}{\partial x}f_{2}(x)$: F(x) = (f_1(x),f_2(x),[f_1,f_2](x)) = (cosx30 sinx3sinx30 - cosx30 1 0 ).
-This matrix is nonsingular for all $x\in{\mathbb{R}}^{3}$, and its inverse is given by
-${\cal F}^{-1}(x)=\begin{pmatrix}\cos x_{3}&\sin x_{3}&0\\ 0&0&1\\ \sin x_{3}&-\cos x_{3}&0\end{pmatrix}$. (2)
-The nonsingularity of ${\cal F}(x)$ implies that the driftless control-affine system (1) satisfies the Lie Algebra Rank Condition (LARC): span (f_1(x),f_2(x),[f_1,f_2](x)) = R^3 for all   x∈R^3.
-Hence, by the Chow–Rashevsky theorem, system (1) is completely controllable in ${\mathbb{R}}^{3}$.
+For the purpose of control design, we introduce the matrix ${\cal F}(x)$ composed of the vector fields $f_{1}(x)$, $f_{2}(x)$ and their Lie bracket $f_{1},f_{2}](x)=\frac{\partial f_{2}(x)}{\partial x}f_{1}(x)-\frac{\partial f_{1}(x)}{\partial x}f_{2}(x)$, and compute ${\cal F}^{-1}(x)$:
+${\cal F}(x)=(f_{1},f_{2},[f_{1},f_{2}])(x)=\begin{pmatrix}\cos x_{3}&0&\sin x_{3}\\ \sin x_{3}&0&-\cos x_{3}\\ 0&1&0\end{pmatrix},{\cal F}^{-1}(x)=\begin{pmatrix}\cos x_{3}&\sin x_{3}&0\\ 0&0&1\\ \sin x_{3}&-\cos x_{3}&0\end{pmatrix}$.
+(2)
+The nonsingularity of ${\cal F}(x)$ implies that system (1) satisfies the Lie Algebra Rank Condition (LARC): ${\rm span}(f_{1}(x),f_{2}(x),[f_{1},f_{2}](x))={\mathbb{R}}^{3}$ for all $x\in{\mathbb{R}}^{3}$.
+Hence, by the Chow–Rashevsky theorem, the driftless control-affine system (1) is completely controllable in ${\mathbb{R}}^{3}$.
 For a given real-valued potential function $V\in C^{2}({\mathbb{R}}^{3})$, the approximate motion of system (1) in the direction of the negative gradient of $V(x)$ is generated by control functions $u^{\omega}(x,t)=(u_{1}^{\omega}(x,t),u_{2}^{\omega}(x,t))^{\top}$ of the following form [10]:
@@
 A comprehensive analysis of this general problem is beyond the scope of the present paper.
-Instead, we restrict our attention to estimating $J_{X}[V]$ within some positive definite quadratic forms.
+Instead, we estimate $J_{X}[V]$ using a selected family of positive definite quadratic forms.
 $c_{1}$ $c_{2}$ $c_{3}$ $J_{X}[V]$
@@
 where $\alpha$ is a parameter.
-In particular, computing the cost (9) for $V_{\alpha}$ with $\alpha=2,4,10$, we obtain: J_X[V_2]≈0.1403,  J_X[V_4]≈0.0962,  J_X[V_10]≈0.0906.
+In particular, computing the cost (9) for $V_{\alpha}$ with $\alpha=2,4,10$, we obtain: $J_{X}[V_{2}]\approx 0.1403,\;J_{X}[V_{4}]\approx 0.0962,\;J_{X}[V_{10}]\approx 0.0906$.
 The obtained sequence of decreasing cost values justifies the use of the Lyapunov function candidates (13) for further experiments.
-It should be emphasized that quadratic forms cannot serve as control-Lyapunov functions for systems of the form (1) due to Brockett’s necessary stabilizability condition.
-Indeed, Brockett’s condition implies that the considered class of nonholonomic systems does not admit any continuously differentiable control-Lyapunov function.
+It should be emphasized that quadratic forms cannot serve as control-Lyapunov functions for systems of the form (1) due to Brockett’s necessary stabilizability condition [6, 18].
+Indeed, Brockett’s condition and Artstein’s theorem [4] imply that the considered class of nonholonomic systems does not admit any continuously differentiable control-Lyapunov function.
 4 Experimental Implementation
@@
 Robotic Platform.
-The experiments were conducted using a TurtleBot3 Burger (TB3) – an affordable, differential-drive mobile robot (see Figure 1, left) [1].
+The experiments were conducted using a TurtleBot3 Burger (TB3) – an affordable differential-drive mobile robot (see Figure 1, left) [1].
 The robot’s control and communication were handled by its onboard Raspberry Pi 4B, running Ubuntu Server 22.04.5 LTS (64-bit) and ROS 2 Humble.
-Mechanically, the TB3 features two off-centered, independently actuated wheels and a caster ball, thus the robot is not perfectly unicycle.
-However, a differential-drive robot can be modeled as unicycle by mapping the velocities of the left and right wheel $v_{l}$ and $v_{r}$ (linear velocities at the wheel edge) to the unicycle velocities $u_{1}$ and $u_{2}$ mentioned in (1).
-Let $d$ be the distance between the two wheels of the differential drive robot, and assume the robot’s center of rotation is at the midpoint between the wheels.
+Mechanically, the TB3 features two off-centered, independently actuated wheels and a caster ball; thus, the robot is not a perfect unicycle.
+However, a differential-drive robot can be modeled as a unicycle by mapping the velocities of the left and right wheels, $v_{l}$ and $v_{r}$ (linear velocities at the wheel edge), to the unicycle velocities $u_{1}$ and $u_{2}$ introduced in (1).
+Let $d$ be the distance between the two wheels of the differential drive robot, and assume that the robot’s center of rotation is at the midpoint between the wheels.
 Then, the unicycle’s translational and angular velocities, respectively $u_{1}$ and $u_{2}$, are given by $u_{1}=\frac{v_{r}+v_{l}}{2},\quad u_{2}=\frac{v_{r}-v_{l}}{d}.$ The inverse of this transformation is handled automatically by the diff_drive_controller of ROS 2.
-It accepts the unicycle-style Twist commands on /cmd_vel and computes the corresponding wheel velocities.
+It accepts unicycle-style Twist commands on /cmd_vel and computes the corresponding wheel velocities.
 The robot’s motion is constrained by its kinematics with the maximum translational velocity of 0.22 m/s and the maximum angular velocity of 2.84 rad/s.
 No onboard sensors were used for state estimation or control, but four reflective markers were mounted for the motion capture system.
+In practice, the TurtleBot3 Burger exhibits limitations beyond velocity constraints, including motor dead zones and nonlinearities.
+Although no encoder or motor velocity feedback was used to compensate for these effects, the experimental results show that the proposed controller still drives the system close to the desired state in a stable manner.
+Incorporating additional feedback into the control law could further improve the performance in more demanding scenarios and is left for future work.
 Motion Capture System.
-Robot pose information was obtained using an OptiTrack motion capture system, consisting of eight infrared cameras and the Motive tracking software.
-As previously stated, four reflective markers were attached to the top of the TB3, allowing the estimation of its x- and y-position as well as the yaw orientation.
-The Motive software typically computes the center point of these markers and uses this as a reference pose.
-However, this point did not align precisely with the robot’s pivot point.
-To ensure accurate control, the reference pose was adjusted manually to match the true center of rotation (see Figure 1, right).
-We used a dedicated wireless access point to connect to the OptiTrack system via LAN, ensuring the TB3 and a remote laptop used for robot control could communicate in a known low-latency wireless network.
+Robot pose information is obtained using an OptiTrack motion capture system, consisting of eight infrared cameras and the Motive tracking software.
+As previously stated, four reflective markers are attached to the top of the TB3, allowing the estimation of its $x$- and $y$-position as well as its yaw orientation.
+The Motive software typically computes the center point of these markers and uses it as a reference pose.
+However, this point does not align precisely with the robot’s pivot point.
+To ensure accurate control, the reference pose is adjusted manually to match the true center of rotation (see Figure 1, right).
+We use a dedicated wireless access point to connect to the OptiTrack system via LAN, ensuring that the TB3 and a remote laptop used for robot control can communicate over a known, low-latency wireless network.
+The robot moves on a standard office carpet floor, which minimizes slip. While the performance may vary on other surfaces, the conducted experiments demonstrate the feasibility of the controller.
+Another important note is that the experiments rely on the OptiTrack system for high-precision position feedback.
+In environments without such a system, odometry-based approaches (i.e., relying only on IMU sensors and motor encoders) would be significantly more challenging.
 Figure 1: Left: TurtleBot3 Burger; Right: TB3 marker configuration
 Software Architecture.
-The control logic for both simulation and real-world experiments was implemented as a single ROS 2 node using the rclpy Python library.
+The control logic for both simulation and real-world experiments is implemented as a single ROS 2 node using the rclpy Python library.
 In real-world experiments, the controller subscribes to the robot pose data provided by the OptiTrack system via the NatNet protocol, whereas in simulations it subscribes to the standard /odom ROS 2 topic published by Gazebo.
@@
 This allows us to use one and the same controller logic in different domains.
-Control commands are recomputed and sent as fast as possible which turned out to be every 0.5 milliseconds.
-Experimental data and controller parameters were recorded automatically for each trial.
-Plots were generated using Python’s matplotlib library, and all data were exported in .csv format for further analysis.
+Although the physical dynamics of the robot are modeled in continuous time, a digital implementation of both controllers (6) and (7) cannot be entirely continuous, as delays are introduced by data retrieval, computation, and data transmission.
+The control commands for the TB3 are generated on a PC and sent to the robot as quickly as possible to implement strategy (7).
+Timestamp measurements indicate that a new control command is issued approximately every 0.5 ms, providing a high-frequency approximation of continuous-time behavior.
+Therefore, for notational convenience, we still refer to this implementation of controller (7) as continuous-time in the sequel.
+Experimental data and controller parameters are recorded automatically for each trial.
+Plots were generated using Python’s matplotlib library, and all data was exported in .csv format for further analysis.
+Figure 2: Experiment 1 – Parameter set $(P1)$: continuous-time vs. sampling.
 Simulation Environment.
-Simulation experiments were conducted using the predefined TB3 Gazebo environment provided in the Robotis e-Manual [2].
-An empty world was used to replicate a simple, obstacle-free arena.
-The same TB3 model (Burger) as in real world was used to ensure consistency in dimensions, kinematics and dynamics.
+Simulation experiments are conducted using the predefined TB3 Gazebo environment provided in the Robotis e-Manual [2].
+An empty world is used to replicate a simple, obstacle-free arena.
+The same TB3 model (Burger) as in real world is used to ensure consistency in dimensions, kinematics and dynamics.
 The model incorporates basic rigid body physics including link masses and joint properties allowing the robot to respond realistically to velocity commands although small inaccuracies in terms of friction and inertia compared to the real-world robot may exist.
 Due to the software architecture the same Python controller could be used.
-Figure 2: Experiment 1 – Parameter set $(P1)$: continuous-time vs. sampling.
 4.2 Experimental Results
@@
 The robot stops automatically if it is in the vicinity of that goal point to obtain comparable convergence time results.
+More precisely, an orientation angle tolerance of $\Delta x_{3}=0.05~rad$ was chosen for all parameter sets (see (P1), (P2), (P3), (P4)).
+While for (P1) and (P2) we chose a target distance tolerance of $\sqrt{x_{1}^{2}+x_{2}^{2}}\leq 0.005~m$, for (P3) and (P4) we chose a target distance tolerance of $\sqrt{x_{1}^{2}+x_{2}^{2}}\leq 0.03~m$ due to longer convergence times.
 The translational and angular velocities are governed by the control law (3), (5) with the frequency parameter $\omega=2\pi$ (i.e., $\varepsilon=1$) and the gain $\gamma=0.05$.
@@
 This choice avoids the need to further reduce $\omega$ and $\gamma$ while still satisfying the input constraints.
-For the experimental validation, we use the Lyapunov function candidate (13) and controls of the form (3), (5) with one of the following sets of parameters:
-$\alpha$ $=1,\;k_{1}=0.5,\;k_{2}=8$ (P1)
-$\alpha$ $=1,\;k_{1}={1}/{\sqrt{2}},\;k_{2}=4\sqrt{2}$ (P2)
-$\alpha$ $=4,\;k_{1}=0.5,\;k_{2}=8$ (P3)
-$\alpha$ $=10,\;k_{1}=0.5,\;k_{2}=8$ (P4)
+For the experimental validation, we use the Lyapunov function candidate (13) and controls of the form (3), (5) with one of the following sets of parameters: $\alpha$ $=1,\;k_{1}=0.5,\;k_{2}=8,$ (P1) $\alpha$ $=1,\;k_{1}=\tfrac{1}{\sqrt{2}},\;k_{2}=4\sqrt{2},$ (P2) $\alpha$ $=4,\;k_{1}=0.5,\;k_{2}=8,$ (P3) $\alpha$ $=10,\;k_{1}=0.5,\;k_{2}=8.$ (P4)
 In Experiments 1 and 2, we examine two different parameter selections $(P1)$ and $(P2)$ with the standard sum-of-squares Lyapunov function candidate $V_{1}(x)=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}$.
@@
 Comparing the continuous-time and sampling approach, we can see that they generate clearly distinct trajectories.
-The continuous time approach follows a relatively direct and steep path which can especially seen from the parameter set $(P1)$, where the robot moves straight $\pi/4$ to the goal without chattering from about $1/3$ of the distance.
+The continuous time approach follows a relatively direct and steep path which can especially be seen from the parameter set $(P1)$, where the robot moves straight $\pi/4$ to the goal without chattering from about $1/3$ of the distance.
 The sampling approach produces a more curved and more chattering but faster path.
@@
 It is remarkable to note that although the theoretical stability proof is valid for sufficiently small values of the parameter $\varepsilon$ (and therefore for sufficiently large control frequency $\omega$), our experiments demonstrate the applicability of these controllers at moderate frequencies with $\varepsilon=1$.
-The choice of control coefficients $k_{1}$ and $k_{2}$ is justified in the present work based on control constraints for differential-drive robots combined with gradient-based control formulas under a realistic choice of a Lyapunov function candidate.
+The choice of control coefficients $k_{1}$ and $k_{2}$ is justified in the present work based on control constraints for differential-drive robots combined with gradient-based controls under a realistic Lyapunov function candidate.
 Second, through the formulation of a suitable optimization problem, an appropriate class of Lyapunov candidates has been studied.
@@
 We leave this open theoretical question for future investigation.
-Another promising direction concerns experimental research aimed at achieving guaranteed finite-time practical stability.
-In this context, by exploiting non-differentiable Lyapunov function candidates, we expect to characterize the time required to reach a prescribed neighborhood of the target point using gradient-flow-type systems.
-This experimental study is left for future development, possibly involving non-Lipschitz generating potential functions.
-Acknowledgement
-Alexander Zuyev gratefully acknowledges the funding by the European Regional Development Fund (ERDF) within the programme Research and Innovation — Grant Number ZS/2023/12/182138.
+Another promising direction concerns experimental research aimed at achieving guaranteed-time practical stabilization to reach a prescribed neighborhood of the target point, possibly exploiting non-differentiable Lyapunov function candidates.